# VSI: Visual–Subtitle Integration for Keyframe Selection to Enhance Long Video Understanding

### v2

@@ Introduction
LongVideoBench Video-MME Model and Size Searching Modality Frame Video Length Model and Size Searching Modality Frame Video Length Long Medium Long Medium 900-3600s 180-600s 30-60min 4-15min GPT-4o (Wu et al. 2024) N/A 4 46.81 48.79 GPT-4o N/A 4 54.75 55.06 GPT-4o + Tstar Unimodal 4 45.39 50.49 GPT-4o + Tstar Unimodal 4 54.64 59.33 GPT-4o + VSLS Unimodal 4 47.52 51.47 GPT-4o + VSLS Unimodal 4 56.76 60.00 GPT-4o + VSI (ours) Multimodal 4 50.09 51.94 GPT-4o + VSI (ours) Multimodal 4 54.75 58.20 GPT-4o (Wu et al. 2024) N/A 8 46.99 50.49 GPT-4o N/A 8 54.53 61.24 GPT-4o + Tstar Unimodal 8 47.34 52.94 GPT-4o + Tstar Unimodal 8 54.97 59.66 GPT-4o + VSLS Unimodal 8 49.11 52.21 GPT-4o + VSLS Unimodal 8 55.64 61.12 GPT-4o + VSI (ours) Multimodal 8 49.20 56.07 GPT-4o + VSI (ours) Multimodal 8 55.75 59.53 PLLaVA-34B+Sub N/A 32 49.10 50.80 GPT-4V+Sub N/A 10 56.9 59.7 LLaVA-Next-Video-34B+Sub N/A 128 45.90 48.70 Claude-3.5-Sonnet+Sub N/A 20 54.70 60.10 GPT-4o-mini+Sub N/A 250 53.40 56.90 LLaVA-OneVision-78B+Sub N/A 32 60.00 62.20 Qwen2-VL-7B+Sub N/A 256 52.50 56.70 TimeMarker-8B+Sub N/A 128 51.90 60.70 GPT-4o(0513)+Sub N/A 256 61.60 66.70 VideoChat-Flash-7B+Sub N/A 512 63.30 63.80
Table 2: Downstream QA task evaluation results on LongVideoBench and Video-MME. Frame indicates the number of frames used in performing VideoQA tasks. Meanwhile, below the main table, we add the performance of the current common LLMs (including the model size and frames used) on VideoQA tasks as a reference for comparison. The data is from the official websites of two datasets (for consistency with our 2-decimal precision, single-decimal values are shown with a trailing zero).
To conclusively demonstrate the superiority of VSI, we conducted downstream VideoQA performance tests on both LongVideoBench and Video-MME datasets. As shown in Table 2, videos are grouped by length into Medium, and Long (the two lengths of video are defined differently in the two datasets, as indicated in Table 2.).
In terms of keyframe search accuracy, as shown in Fig  3, regardless of the premise of taking top-4 frames or top-8 frames, VSI achieves SOTA on the LongVideoBench full dataset, with 29.81% and 38.40% keyframe search accuracy, respectively, which is an improvement of 1.10% and 0.33% compared with the second-best unimodal methods. It shows that our multimodal search framework is not only adaptable to keyframe search for text-related perception tasks, but also effective in keyframe search for other types of tasks, which reflects the generalization of VSI.
In terms of the correct rate of downstream QA tasks, for the LongVideoBench dataset, VSI achieved SOTA performance in both the medium video QA task as well as the long video QA task. When inputting only top-4 frames to the GPT4o for QA, our methods achieve 51.94% and 50.09% correct rates in the medium video QA task and long video QA task, respectively, which are higher than the baseline method by 3.15% and 3.28%, and also improve by 0.47% and 2.57% compared to the second-best unimodal method, respectively. The excellent performance of VSI on the LongVideoBench dataset illustrates the correlation between the correct selection of keyframes and the correct answers of the downstream tasks on the one hand, and the robustness of the VSI framework on the other. For Video-MME, since this dataset does not have annotations for ground truth frames, we are currently unable to test keyframe search accuracy on this dataset. However, VSI also performed well in downstream video-QA tasks. Based on inputting the top-8 frames, our method achieves SOTA performance on the long Video-QA task. Additionally, when inputting the top-4 frames, we improve GPT-4o’s answer accuracy rate on the Medium Video-QA task from 55.06% to 58.20%.
Analysis
Method Training Required Searching Efficiency Frame Acc $\uparrow$ Video Search Text Encoding Iteration $\downarrow$ Tstar-Attention(Ye et al. 2025) Training-Based N/A N/A N/A N/A N/A Tstar-Detector(Ye et al. 2025) Training-Free YOLO-World-110M N/A 43 4 27.74 VSLS-Detector(Guo et al. 2025) Training-Free YOLO-World-110M N/A 49 4 28.71 VSI-Detector(ours) Training-Free YOLO-World-110M all-mpnet-base-v2 26 4 29.81 Tstar-Attention(Ye et al. 2025) Training-Based N/A N/A N/A N/A N/A Tstar-Detector(Ye et al. 2025) Training-Free YOLO-World-110M N/A 43 8 36.84 VSLS-Detector(Guo et al. 2025) Training-Free YOLO-World-110M N/A 49 8 38.07 VSI-Detector(ours) Training-Free YOLO-World-110M all-mpnet-base-v2 26 8 38.40
@@
Table 3: Evaluation of computational performance metrics across the LongVideoBench benchmark. Training Required shows whether the method needs to be trained, Video Search indicates the model used by the method for video modality processing, and Text Encoding indicates the model used by the method for text modality processing. Iteration represents the average number of iterations needed for each keyframe search.
Computational Efficiency Metrics
Prior research (Fan et al. 2024; Park et al. 2024; Wang et al. 2024b, c; Wu and Xie 2023) has predominantly focused on enhancing domain-specific accuracy metrics, often overlooking the computational overhead of temporal search processes. To comprehensively assess this aspect, our evaluation scheme introduces two quantitative measures:
1) Computational Complexity measured in FLOPs:
$\text{FLOPs}=\sum_{l=1}^{L}(2\cdot C_{l}^{in}\cdot C_{l}^{out}\cdot K_{l}^{2}\cdot H_{l}\cdot W_{l}+C_{l}^{out}\cdot H_{l}\cdot W_{l}),$ (16)
where $L$ denotes total layers, $C_{l}$ represents channels, $K_{l}$ the kernel size, and $H_{l}\times W_{l}$ the spatial dimensions at layer $l$.
Time Complexity
The proposed multimodal framework extends the original approach by incorporating textual semantic analysis alongside visual processing. The enhanced method maintains the original two-stage architecture while introducing a parallel text processing pipeline. Our complexity analysis focuses on the integrated workflow:
Initial Phase: The system simultaneously processes visual content through VLMs and analyzes subtitle text using a sentence transformer (e.g., all-mpnet-base-v2). This dual processing generates:1)Visual semantic set $\mathcal{S}_{v}$ (target/cue objects and visual relations) ;2)Textual semantic set $\mathcal{S}_{t}$ (temporal anchors and textual cues)
Search Phase: The algorithm now operates with an augmented semantic set $\mathcal{S}=\mathcal{S}_{v}\cup\mathcal{S}_{t}$, implementing a hybrid scoring mechanism:
$h^{\prime}(f,\mathcal{S})=h_{v}(f,\mathcal{S}_{v})+\alpha_{t}h_{t}(f,\mathcal{S}_{t}),$ (17)
where $\alpha_{t}$ controls text-visual balance.
@@
Complexity Analysis: The multimodal search retains the original $\mathcal{O}(|\mathcal{S}|k\log n)$ YOLO-World detection complexity, with three key additions: 1) Text Processing Overhead: Subtitle analysis introduces $\mathcal{O}(m)$ complexity for $m$ text segments, which is negligible compared to visual processing; 2) Search Convergence: Text-guided sampling empirically reduces the number of iterations by 38%; 3) Fusion Cost: Cross-modal scoring adds a constant $\mathcal{O}(1)$ overhead per frame.
Bottleneck Preservation: YOLO-World detection remains the computational bottleneck. Our implementation shows only 9.7% increased inference time versus the visual-only baseline, demonstrating efficient multimodal integration.
Evaluation of Search Framework Efficiency
As shown in Table 3, based on the complementarity of textual and visual information, the frame scores provided by the Subtitle Match stream in VSI can assist in updating the frame sampling probability in the Video Search stream. Therefore, the object detection model (YOLO-World-110M used in this paper) can find all objects more quickly. Therefore, on the LongVideoBench dataset, compared to single-modal methods, the VSI achieve higher keyframe search accuracy. While the next-best method requires an average of 49 iterations, our method only needs an average of 26 iterations to complete a search process, demonstrating the efficiency and effectiveness of our method.
In order to serve VSI lightweight and efficient design purpose, we have also designed in the model selection: 1) In the video search model, we chose YOLO-World-110M (Cheng et al. 2024a). This is a real-time open vocabulary target detection model developed by Tencent AI Lab. YOLO-World-110M realizes real-time target detection with open vocabulary, and can locate objects with zero samples through natural language descriptions, its lightweight architecture combined with the heavy parameterization convolution technique also meets our lightweight needs; 2) In text coding model we chose all-mpnet-base-v2 (Reimers and Gurevych 2020), which is a sentence embedding model based on the pre-trained model microsoft/mpnet-base (Song et al. 2020). It demonstrates excellent robustness in standard semantic search scenarios. Its optimized contrastive learning mechanism calculates cosine similarity between similar texts more accurately and is good at handling text matching between sentences.
Conclusion
This paper presented Visual-Subtitle Integration (VSI), a dual-stream keyframe retrieval framework that integrates subtitle-based semantic matching with visual object detection to guide frame sampling for long-video understanding. By aligning the query with both time-stamped captions and detected objects, VSI updates the sampling distribution toward frames that are more relevant for reasoning, improving localization while keeping computation practical. Experiments show that VSI achieves SOTA performance on LongVideoBench both in terms of keyframe search accuracy and downstream Video-QA tasks, and the method shows consistent performance across different backbone architectures. Being training-free and plug-and-play, VSI can be easily integrated into existing video-language pipelines with minimal modifications.

### v3

@@ 3 Exeriments
LongVideoBench Video-MME Model Searching Modality Frame Video Length Model Searching Modality Frame Video Length Long Medium Long Medium GPT-4O [hurst2024gpt] N/A 8 47.0 50.5 GPT-4O N/A 8 54.5 61.2 GPT-4O + Tstar [ye2025rethinking] Unimodal 8 47.3 52.9 GPT-4O + Tstar Unimodal 8 55.0 59.7 GPT-4O + VSLS [vsls] Unimodal 8 49.1 52.2 GPT-4O + VSLS Unimodal 8 55.6 61.1 GPT-4O + VSI (ours) Multimodal 8 49.2 56.1 GPT-4O + VSI (ours) Multimodal 8 55.8 59.5 LLaVA-Video-7B-Qwen2 [36] N/A 8 39.0 42.7 LLaVA-Video-7B-Qwen2 N/A 8 38.0 39.7 LLaVA-Video-7B-Qwen2+Tstar Unimodal 8 40.8 44.1 LLaVA-Video-7B-Qwen2+Tstar Unimodal 8 37.5 40.4 LLaVA-Video-7B-Qwen2+VSLS Unimodal 8 38.7 42.4 LLaVA-Video-7B-Qwen2+VSLS Unimodal 8 38.5 38.5 LLaVA-Video-7B-Qwen2+VSI (ours) Multimodal 8 40.6 44.1 LLaVA-Video-7B-Qwen2+VSI (ours) Multimodal 8 38.8 42.2 QWEN2.5-VL-7B-INSTRUCT [qwen] N/A 8 38.7 43.5 QWEN2.5-VL-7B-INSTRUCT N/A 8 38.0 47.3 QWEN2.5-VL-7B-INSTRUCT+Tstar Unimodal 8 39.2 43.9 QWEN2.5-VL-7B-INSTRUCT+Tstar Unimodal 8 40.3 50.0 QWEN2.5-VL-7B-INSTRUCT+VSLS Unimodal 8 42.4 43.9 QWEN2.5-VL-7B-INSTRUCT+VSLS Unimodal 8 43.2 49.6 QWEN2.5-VL-7B-INSTRUCT+VSI (ours) Multimodal 8 39.9 45.1 QWEN2.5-VL-7B-INSTRUCT+VSI (ours) Multimodal 8 40.8 48.5 GPT-4O N/A 32 48.2 51.9 GPT-4O N/A 32 55.2 61.0 GPT-4O + Tstar Unimodal 32 46.6 52.5 GPT-4O + Tstar Unimodal 32 55.2 61.6 GPT-4O + VSLS Unimodal 32 49.1 50.7 GPT-4O + VSLS Unimodal 32 57.5 61.9 GPT-4O + VSI (ours) Multimodal 32 51.2 53.9 GPT-4O + VSI (ours) Multimodal 32 55.3 61.7 LLaVA-Video-7B-Qwen2 N/A 32 39.4 45.3 LLaVA-Video-7B-Qwen2 N/A 32 35.9 36.5 LLaVA-Video-7B-Qwen2+Tstar Unimodal 32 39.2 43.9 LLaVA-Video-7B-Qwen2+Tstar Unimodal 32 35.8 39.6 LLaVA-Video-7B-Qwen2+VSLS Unimodal 32 42.4 43.9 LLaVA-Video-7B-Qwen2+VSLS Unimodal 32 36.9 39.0 LLaVA-Video-7B-Qwen2+VSI (ours) Multimodal 32 39.9 45.3 LLaVA-Video-7B-Qwen2+VSI (ours) Multimodal 32 38.2 42.2 QWEN2.5-VL-7B-INSTRUCT N/A 32 30.3 35.0 QWEN2.5-VL-7B-INSTRUCT N/A 32 37.5 39.9 QWEN2.5-VL-7B-INSTRUCT+Tstar Unimodal 32 38.7 42.9 QWEN2.5-VL-7B-INSTRUCT+Tstar Unimodal 32 34.9 45.6 QWEN2.5-VL-7B-INSTRUCT+VSLS Unimodal 32 38.1 43.1 QWEN2.5-VL-7B-INSTRUCT+VSLS Unimodal 32 37.9 50.0 QWEN2.5-VL-7B-INSTRUCT+VSI (ours) Multimodal 32 37.1 44.4 QWEN2.5-VL-7B-INSTRUCT+VSI (ours) Multimodal 32 39.3 45.8
3.1 Benchmark Datasets
We comprehensively evaluated the proposed VSI framework on two pivotal benchmarks: LongVideoBench [lvb] for assessing long-context video-language understanding capabilities, and Video-MME [videomme] as the first integrated benchmark for multimodal video analysis focused on QA evaluation. Leveraging ground-truth key frame annotations in LongVideoBench enables rigorous validation of key frame selection accuracy, while both datasets collectively support quantitative analysis of VideoQA performance, with detailed dataset statistics provided in the Appendix.
3.2 Search effectiveness and efficiency
For the quantitative evaluation of search performance, we adopt keyframe search accuracy as the core metric. Given that the LongVideoBench dataset provides authoritative ground truth annotations for key frames, we define this accuracy mathematically as follows:
Let the ground truth keyframe index of a video be $g$, and the set of top-k frames selected based on score distribution be $F=\{f_{1},f_{2},\ldots,f_{k}\}$ where $f_{i}$ denotes the index of the $i$-th frame in the top-k result. A search is deemed valid if there exists at least one frame in $F$ such that the frame index deviation from $g$ does not exceed 200 frames, i.e.,
$\exists f\in F\quad\text{s.t.}\quad|f-g|\leq 200$ (16)
Let $N_{\text{valid}}$ denote the number of videos with valid searches, and $N_{\text{total}}$ denote the total number of searched videos. The keyframe search accuracy is then defined as the proportion of validly searched videos relative to the total number of searched videos:
$\text{Accuracy}=\frac{N_{\text{valid}}}{N_{\text{total}}}$ (17)
This definition ensures that the retrieved frames still retain the core semantics of the ground truth (GT) keyframe. Specifically, these proximate frames adequately enable subsequent VLMs to comprehend critical video content, thus avoiding the loss of utility of retrieval results for downstream tasks caused by excessive frame deviation.
To conduct a systematic quantitative evaluation of search efficiency, this paper employs three core metrics: TFLOPs, Iteration, and Latency(sec). Specifically, TFLOPs (Trillions of Floating-Point Operations Per Second) quantifies the computational complexity of the search process; Iteration denotes the average number of iterations required to complete a single keyframe search task; and Latency represents the average time elapsed from the initiation to the completion of a single keyframe search task.
In this study, Tstar and VSLS—two approaches that also employ a keyframe search architecture—are selected as baseline models. As shown in Table 2, under the top-64 frame setting, the VSI method achieves state-of-the-art (SOTA) performance in keyframe search accuracy on the full LongVideoBench dataset, as well as on its image-only subset and text-only subset individually. Specifically, compared to the second-best performing methods, VSI demonstrates accuracy improvements of 3.66%, 1.22%, and 10.81% across these three evaluation scenarios, respectively. The improvement observed on the text-only subset aligns with intuitive expectations, indicating that the proposed method effectively leverages textual modality information to enhance search accuracy in this context. Notably, VSI also outperforms unimodal keyframe search methods on the image-only subset alone, suggesting that the subtitle matching stream provides complementary textual information that supports the video search stream, thereby significantly boosting keyframe search accuracy. This outcome underscores the robustness and adaptability of the VSI framework across diverse query scenarios. Furthermore, VSI achieves these performance gains without incurring significant increases in computational complexity, as measured by TFLOPs, average iteration count, or search latency, thus maintaining an efficient balance between retrieval effectiveness and efficiency.
3.3 Downstream VideoQA Performance
@@
Table 2: Evaluation of computational performance metrics across the LongVideoBench benchmark. Training Required shows whether the method needs to be trained, Video Search indicates the model used by the method for video modality processing, and Text Encoding indicates the model used by the method for text modality processing.
Searching Efficiency Method Training Required Video Search Text Encoding TFLOPs Iteration Latency(sec) Frame Acc $\uparrow$ ALL T*-DETECTOR [ye2025rethinking] Training Free YOLO-World-110M N/A 31.7 24.76 28.96 64 67.58 VSLS-DETECTOR [vsls] Training Free YOLO-World-110M N/A 33.3 24.49 33.26 64 70.23 VSI-DETECTOR(ours) Training Free YOLO-World-110M ALL-MPNET-BASE-V2 36.8 26.34 31.71 64 73.89 IMAGE-ONLY T*-DETECTOR Training Free YOLO-World-110M N/A 31.7 23.90 25.4 64 70.34 VSLS-DETECTOR Training Free YOLO-World-110M N/A 33.3 24.04 28.13 64 68.62 VSI-DETECTOR(ours) Training Free YOLO-World-110M ALL-MPNET-BASE-V2 36.8 25.35 30.92 64 71.56 TEXT-ONLY T*-DETECTOR Training Free YOLO-World-110M N/A 31.7 26.80 27.99 64 66.36 VSLS-DETECTOR Training Free YOLO-World-110M N/A 33.3 26.39 28.77 64 66.36 VSI-DETECTOR(ours) Training Free YOLO-World-110M ALL-MPNET-BASE-V2 36.8 28.37 34.44 64 77.17
To conclusively demonstrate the superiority of VSI, we conducted downstream VideoQA performance tests on both LongVideoBench (LVB) and Video-MME (VMME) datasets. We focus on the performance on Medium(LVB:3-10min,VMME:4-15min) and Long(LVB:10-60min,VMME:30-60min) videos.
For the full-scale VideoQA task, we selected GPT-4o, LLaVA-Video-7B-Qwen2, and Qwen2.5-VL-7B-instruct as foundation models. As shown in Table 1, experiments were conducted with two input frame configurations: 8 frames and 32 frames. Experimental results demonstrate that for medium-length video tasks on LVB, VSI achieves state-of-the-art (SOTA) performance under both input frame configurations. Furthermore, across all experimental settings, VSI consistently exhibits significant performance improvements compared to baseline methods employing uniform sampling strategies, fully validating its effectiveness in medium-to-long video QA scenarios.
3.4 Visual Subtitle Integration Performance
To substantiate our method’s efficacy in leveraging textually relevant information, we evaluated keyframe search accuracy on the text-related perception subset of LongVideoBench. Specifically, we define the Text-Related Perception Subset as encompassing three task types: Text-referred Object Attribute (T2A), Text-referred Event (T2E), and Text-referred Object (T2O). Given that key frame localization in these tasks exhibits strong dependence on subtitle information, this subset provides direct empirical validation of our method’s superiority over unimodal approaches. Detailed dataset specifications are provided in the appendix. As shown in Table 3, VSI reaches 40% key frame search accuracy with only top-4 frames, a 20.35% gain over the second-best unimodal method. For downstream VideoQA’s medium and long video tasks, it achieves 63.79% and 68.48% accuracy, improving by 10.34% and 14.72% over the second-best method.
Notably, the LongVideoBench benchmark includes a subtitle-based baseline method: this approach uniformly samples k video frames(e.g., 8) and feeds their corresponding subtitles as prompts to GPT-4o (denoted GPT-4o+Sub). This baseline reaches 58.06% accuracy on long video QA tasks, exceeding caption-free methods by 4.30% and verifying the necessity of subtitles. In contrast, VSI using top-8 frames achieve 69.67% accuracy on the same task, outperforming this subtitle baseline by 11.51%.
Table 3: Keyframe Search Accuracy and Downstream QA Accuracy on LongVideoBench-Text-Related Perception Subsets.
Method Frame Acc $\uparrow$ LongVideoBench Medium Long GPT4o [lvb] 4 N/A 46.55 46.24 GPT4o+Sub [lvb] 4 N/A 44.83 45.16 GPT4o+Tstar [ye2025rethinking] 4 19.65 53.45 53.76 GPT4o+VSLS [vsls] 4 18.50 50.00 52.69 GPT4o+VSI (ours) 4 40.00 63.79 68.48 GPT4o [lvb] 8 N/A 48.28 53.76 GPT4o+Sub [lvb] 8 N/A 46.55 58.06 GPT4o+Tstar [ye2025rethinking] 8 29.48 48.28 58.06 GPT4o+VSLS [vsls] 8 27.17 48.28 55.91 GPT4o+VSI (ours) 8 45.00 62.07 69.57
4 Analysis
4.1 Video Search and Subtitle Match
To further quantify the independent contributions of the Video Search and Subtitle Match branches, we also tested and reported the performance of the standalone Video Search and Subtitle Match branches on the image-only and text-only subsets of the LongVideoBench dataset. As shown in Table 4, keyframe search accuracy results indicate that across all experimental settings on the image-only subset, the Video Search branch consistently achieves higher search accuracy than the Subtitle Match branch. Conversely, on the text-only subset, the Subtitle Match branch demonstrates significantly superior search performance compared to the Video Search branch. This branch also exhibits excellent adaptability in the downstream VideoQA task.
From the perspective of dual-branch interaction, during the keyframe search phase, the Video Search branch undertakes the core search task and serves as the primary module for completing keyframe retrieval. The Subtitle Match branch plays a supporting role, primarily injecting text-based semantic information that complements the image modality into the Video Search branch. Assisting this branch in dynamically updating its sampling probability distribution enables more precise targeting of video segments containing core semantic information. In the final keyframe selection phase, since keyframe selection relies on the combined score distribution output by both branches, the contribution weight of each branch to this distribution is, to a certain extent, determined by the relevance match between the current VideoQA query and either the image modality or the text modality.
As shown in Figure 3, the video timestamp corresponding to the subtitle with the highest semantic similarity matched by the Subtitle Match branch is 127.79s–127.80s. Subsequently, through the Score Fusion mechanism, this key textual semantic information is injected into the sampling probability distribution of the Video Search branch. This process enables the branch to focus more precisely on video frame regions containing core information relevant to the query during iterative searches, fully demonstrating the guiding role of text-modal information in visual search.
Table 4: Ablation Study of Video Search branch and Subtitle Match branch. TEXT-ONLY and IMAGE-ONLY refer to tasks within the Text-Referred category on LongVideoBench and all other task types. VideoQA uses GPT4o as the base model.
Method Frame Acc $\uparrow$ LongVideoBench Long Medium IMAGE-ONLY VSI-Subtitle Match 4 20.18 49.56 51.91 VSI-Video Search 4 23.55 48.09 52.67 VSI-Subtitle Match 8 27.68 50.15 54.58 VSI-Video Search 8 34.56 49.27 59.16 TEXT-ONLY VSI-Subtitle Match 4 40.00 68.48 60.34 VSI-Video Search 4 14.37 60.22 53.45 VSI-Subtitle Match 8 45.00 69.57 62.07 VSI-Video Search 8 20.00 58.06 55.17
4.2 Ablation Study of Text Weight
To further quantify the independent utility of the Subtitle Match Branch, we conduct ablation experiments targeting the core parameter Text Weight. This parameter’s primary function is to dynamically adjust the weighting proportion of the Subtitle Match Branch’s output scores during the score fusion stage of keyframe search. As shown in Table 5, experiments on the Text-Related Perception Subsets of the LongVideoBench dataset reveal a significant pattern: As Text Weight progressively increases, both keyframe search accuracy and answer correctness in the downstream VideoQA task show synchronous improvement. This dynamic outcome fully validates that when addressing highly text-related VideoQA problems, the Subtitle Match Branch provides crucial text-modal support. Its presence is indispensable, further confirming the rationality of this design.
Table 5: Keyframe Search Accuracy and Downstream QA Accuracy on LongVideoBench-Text-Related Perception Subsets.
Method Text Weight Frame Acc $\uparrow$ LongVideoBench Medium Long GPT4o+VSI 0.3 4 14.47 45.61 56.99 GPT4o+VSI 0.7 4 18.99 55.17 54.95 GPT4o+VSI 1 4 40.00 63.79 68.48 GPT4o+VSI 0.3 8 20.75 45.61 56.99 GPT4o+VSI 0.7 8 28.48 56.9 60.44 GPT4o+VSI 1 8 45.00 62.07 69.57
4.3 Time Complexity
@@
The proposed multimodal framework extends the unimodal approach by incorporating textual semantic analysis alongside visual processing. The enhanced method maintains the original two-stage architecture while introducing a parallel text processing pipeline. The multimodal search retains the original $\mathcal{O}(|\mathcal{S}|k\log n)$ YOLO-World detection complexity, with two key additions: 1) Text Processing Overhead: Subtitle analysis introduces $\mathcal{O}(m)$ complexity for $m$ text segments, which is negligible compared to visual processing; 2) Fusion Cost: Cross-modal scoring adds a constant $\mathcal{O}(1)$ overhead per frame. YOLO-World detection remains the computational bottleneck. Our implementation shows only 10.5% increased inference time versus the visual-only baseline, demonstrating efficient multimodal integration.
5 Related work
Challenges in Long-Form Video Understanding. Long-form video analysis [xun2025rtv, liu2025javisgpt] poses greater and more persistent challenges than short-video or image tasks [survey], primarily due to inherently complex temporal dynamics and substantial content redundancy [qian2024streaming, zeng2024timesuite, yu2019activityqa, dang2024exploring]. The large number of frames introduces significant computational burdens, substantially increasing memory and latency, making exhaustive frame-wise processing infeasible. Moreover, key semantic events are often temporally sparse and discontinuous, requiring specialized architectures capable of capturing subtle, long-range dependencies [ranasinghe2025understanding, shi2024unlocking, chen2024rextime, weng2024longvlm]. The visual complexity of such videos also introduces noisy, ambiguous, or irrelevant signals, which calls for more effective content distillation mechanisms [84, cheng2024videollama2advancingspatialtemporal, xu2023retrieval, ye2025rethinking, hjx, fmri, d].
Video-Language Modeling Strategies. Recent work using VLMs has primarily focused on three complementary directions: 1) content-aware sampling for more efficient computation [buch2025flexible, yu2025framevoyager, hu2025mllm, tspo, moss], 2) hierarchical reasoning for complex query understanding [wang2024videoagent, min2024morevqa, liao2024align], and 3) information compression to better mitigate context window limitations [tao2025dycoke, liu2025hicom, choudhury2024rlt, chen2025longvila]. Representative strategies include segmentation-retrieval pipelines that divide long videos into clips and employ either learned or heuristic mechanisms to reliably select relevant segments [pan2023retrieving, choudhury2023zero, rohan2025video], and fine-grained frame-level token compression to significantly reduce computational costs [li2024llamavid, chen2024llavolta, song2024less], or adopt memory propagation techniques to better maintain temporal continuity [qian2024streaming, wu2022memvit, liu2024llava]. Recent advances also explore planning systems that decompose complex tasks into sequential perception steps, thereby enabling dynamic and adaptive frame-level information acquisition [wang2024videoagent, liao2024videoinsta].
6 Conclusion
This paper proposes Visual-Subtitle Integration (VSI), a multimodal keyframe search framework, whose core innovation lies in a dual-branch cross-modal collaborative retrieval mechanism combining Video Search and Subtitle Match. Through deep fusion of information from both branches, VSI achieves semantic-level precise localization of keyframes. On the LongVideoBench dataset, VSI demonstrates highly efficient retrieval performance: achieving 73.89% keyframe retrieval accuracy while sampling only 3.2% of video frames on average, enabling VLMs to outperform the uniform sampling baseline in downstream VideoQA performance. Moreover, VSI demonstrates adaptability and robustness in text-related tasks. VSI operates as a plug-and-play solution requiring no additional training or annotation, enabling seamless integration into existing video processing pipelines with outstanding engineering practicality.